OpenStudy (anonymous):

solve by elimination method 9x-9y=27 7y-3x=-11

OpenStudy (anonymous):

time (3) u get : 21y-9x =-33 , -9x+21y=-33 9x-9y=27 -9x+21y=-33 Add : ------------- 12y =-6 y= -6/12 y =-1/2 Sutitution y= -1/2 solve for x 9x-9(-1/2)=27 9x+9/2= 27 9x=27-(9/2) x=(45/2)/9 x=45/18 x=5/2

OpenStudy (anonymous):

nancy thanks but the answer is (2.5,-0.5)

OpenStudy (anonymous):

yes the same -1/2=-0.5 5/2=2.5

OpenStudy (anonymous):

Ty